OpenStudy (anonymous):

A 7500-kg rocket blasts off vertically from the launch pad with a constant upward acceleration of 2.25 m/s2 and feels no appreciable air resistance. When it has reached a height of 525 m, its engines suddenly fail so that the only force acting on it is now gravity. a) What is the maximum height this rocket will reach above the launch pad? b) How much time after engine failure will elapse before the rocket comes crashing down to the launch pad?

OpenStudy (anonymous):

motion equation v^2 = u^2 + 2as u=0 s=525 a=10 v^2 = 2362.5 again using that when rocket reaches highest height v=0 u^2 = 2362.5 a= -10 Solving s = 118.125 total height = 118.125 + 525= 643.125 m b) s=ut + 1\2at^2 s= 643.125 u=0 a=10 solving t= 11.34s

OpenStudy (anonymous):

thanks, we have the same answer but my teacher has a different answer in letter b... her answer is 16.44 but I keep on getting 11point something :(

OpenStudy (anonymous):

btw thanks